There are two plans for joining the steve math club. In the first you pay $250 for an annual membership and $8 per meeting. In the second plan you just pay $13 per meeting. After how many meetings are the plans the same cost? Set up an algebraic equation and solve.
Answer by stanbon(48568)

A team is to design a cover a box to hold a jigsaw puzzle. The design will use 525 square cm. The cardboard is precut to 29 cm x 33 cm. The depth of the box (x) can vary to fit the needs of your cover design. How deep should the box be in order to allow the area of 525 cm for your design?
------------------
Draw the picture of a 29 by 33 rectangle
Cut squares that are x by x out of each corner of the rectangle.
Fold the 4 sides of the rectable up to form the depth of the cover.
--------------------------
The base of the rectangle must be 525 cm^2.
-------------
The dimensions of the base are 29-2x by 33-2x
----------------------------------------------------
Solve (29-2x)(33-2x) = 525
957 -58x - 66x + 4x^2 = 525
4x^2 -124x + 432 = 0
Simplify:
x^2 - 31x + 108 = 0
(x-4)(x-27) = 0
Reasonable solution;
x = 4 cm (depth of the box cover)
=====================================
Cheers,
Stan H.